OpenStudy (anonymous):

What is the correct solution to the equation? 4(x – 3) = 16 A. 1 B. 7 C. 15 D. 24

OpenStudy (anonymous):

4(x-3) = 16 Use the distributive property 4x - 12 = 16 Add 12 to both sides 4x = 28 Divide each side by 4 x = 7 Does that make sense?

OpenStudy (anonymous):

mhmm

OpenStudy (anonymous):

Good :)

OpenStudy (anonymous):

What is the correct solution to this problem? A number divided by 5 is 3 more than –6. What is the number? A. –45 B. –30 C. –15 D. –5

OpenStudy (anonymous):

The answer is C, but if you want a further explanation please make another question.

OpenStudy (anonymous):

k give me a few to right it @KirbyLegs

OpenStudy (anonymous):

Which line shows the first error in the solution? 8x-6=4x-2 (1) 8x-4x-6=4x-4x-2 (2) 4x-6=2 (3) 4x-6+6=2+6 (4) 4x=8 (5) 4x/4= 8/4 (6) x=2 (7) A. line 3 B. line 4 C. line 5 D. line 6 @KirbyLegs

OpenStudy (anonymous):

im ready @KirbyLegs